OpenStudy (anonymous):

A bat was flying for six hours at an average speed of 7 kph. The bat flew at 11 kph during the first two hours of its flight. How fast did the bat fly the last 5 hours?

OpenStudy (calculusxy):

Do you mean how fast the bat flew in the last 4 hours?

OpenStudy (calculusxy):

Because if you are saying "six hours" at the beginning of the question and then ask for 5 hours, it kind of doesn't make much sense to me.

OpenStudy (calculusxy):

But if the question is asking for 4 hours then you would have to do: \[\large \frac{ 11 + 11 + 4x }{ 6 } = 7\] I put 11 twice, or you could simply put 22 to show the first two hours and then the speed for the remaining 4 hours would be unknown, or x. To find average we have to divide the total from the data set by the number of terms in the data set so i divided that by 6. And the average should be 7 (which is what the equation is equal to).

OpenStudy (calculusxy):

And then solve for \(\large x\).

OpenStudy (anonymous):

what did you get for x? bc what I'm getting doesn't make sense.

OpenStudy (calculusxy):

\[\large \frac{ 22 + 4x }{ 6 } = 7 \rightarrow 22 + 4x = 42 \rightarrow 4x = 20 \rightarrow x = 5 \]